the weight ratio of o2 and n2is 1:4 are mixed find ratio of no.of molecules

Grade:12

1 Answers

vikas askiitian expert
13 years ago

number of molecules is directly proportional to moles

N = kn(moles)

let weight of O2 of M then weigth of N2 is 4M ..

moles of O2 = M/32                              (molar mass of O2 is 32)

moles of N2 = 4M/28                             (molar mass of N2 is 28)

 

N02 = k(M/32)             ..............1

Nn2 = k(4M/28)            .............2

 

dividing both the equations

No2/Nn2 = 7/32

 

approve my ans if u like it